# Cryogenic source of atomic tritium for precision spectroscopy and neutrino-mass measurements

### v1

@@ II.4 Scattering of H isotopes on each other and on He
Elastic collisions that do not change the spin state of colliding atoms lead to the recovery of thermal equilibrium and are essential for evaporative cooling. Inelastic rates typically lead to a relaxation to the untrapped states and loss of the atoms. Scattering of atoms with helium vapor is important for the transport of hydrogen into the trapping cell and defines the efficiency of the buffer gas cooling, which is one of the key effects in this proposal.
@@
In the zero temperature limit, the elastic collision cross-section is related to the s-wave scattering length as , where is the two-body correlator defined by quantum statistics and identical properties of colliding particles. 1, 2, 0 for two distinguishable particles, for two indistinguishable bosons, and for two identical fermions accordingly. The value and sign of the scattering length depend on the interaction potential and the electron spin state of the atoms of the pair. The singlet potential supports a large number of bound molecular states and is not relevant for spin-polarized or spinless atoms considered here. The repulsive triplet potential has a weak well because of the Van-der-Waals attraction and may support weakly bound states, or dimers. The shape of the potential is nearly the same for all atomic pairs considered here. Its depth increases with the reduced mass of the atomic pair, and this turns out to be the main parameter in the scattering properties. The dimer bound state occurs for the heaviest pair considered here 4He-4He pair. The pair of T atoms is very close to the binding resonanceElliott and Jones (2025). At the scattering resonance, the scattering length diverges and changes sign, going from to . The scattering length and the collisional cross-section then depend on how far the reduced mass is from the resonance value of Elliott and Jones (2025). We present scattering lengths and cross-sections for elastic collisions of various pairs of H and He in Table 2. Of particular significance for this work is a very large elastic collision cross-section for the heaviest pairs considered, and especially for the T-T collisions.
| D-D | T-T | H-T | T-3He | T-4He | 3He-3He | 3He-4He | 4He-4He | |
| , Å | 3.7 | -42.3 | -0.85 | -7* | -16.9* | -7 | -16.9 | -122.4 |
| , Å2 | 0 |
@@
Processes that occur in inelastic collisions are the spin-exchange and dipolar relaxation. The first class concerns the collisions of atoms in different hyperfine states leading to a rapid relaxation and removal from the trap of the so-called mixed states (c for H/T and and for D). The remaining gas in the pure (doubly polarized) states d (H/T) and (D) are much more stable and may relax only due to the dipole interaction. The dipolar relaxation rates define the stability of the trapped doubly polarized gas. Since the relaxation in two body collisions is a second order process in density, the characteristic decay time is inversely proportional to the gas density with the dipolar relaxation rate constant cm3/s, independent of temperature. The exchange and dipolar relaxation rates for numerous channels for atomic hydrogen were calculated in Refs. Stoof et al. (1988); Elliott and Jones (2025) and experimentally confirmed in the work of the MIT Hess et al. (1987); Doyle (1991) and Amsterdam groups van Roijen et al. (1988). For D, theory predicts two orders of magnitude faster spin-exchange rates, and the dipolar relaxation decreases linearly with temperature as cm3/s. Remarkably, the doubly polarized D gas is becoming more stable at lower temperatures, and its decay time at 1 mK is two orders of magnitude larger than that for H, while the thermalization rate remains fairly largeKoelman et al. (1988). Compared with the potential of all three isotopes for precision spectroscopy in the magnetic traps, deuterium is obviously the best candidate. For tritium, relaxation rate constants data were recently calculated Elliott and Jones (2025). Spin-exchange and dipolar relaxation are much larger than for H, and the rates have a complex dependence on temperature and magnetic field. Under optimal conditions of magnetic fields below 0.1 T and a temperature of 100 mK the dipolar relaxation rate is 5 times larger than for H. Unfortunately, this implies that the fluxes of T in loading the trap should be increased by the same factor.
@@ II.5 Stabilization of H, D, and T in solid molecular matrices
Several methods are known to produce the matrices with embedded atoms. We briefly describe two methods relevant to this work: dissociation with cryogenic rf discharge and electrons resulting from -decay of T2. The first technique is based on an in situ dissociation of the matrix molecules using pulsed rf discharge in the helium vapor above solid films of hydrogens below 1 K introduced by Ahokas et.al. Ahokas et al. (2009). The molecules are split by the impact of the electrons of the discharge having energies of the order of 100 eV. It turned out that a fraction of the atoms is evaporated during the rf pulse, and the H and D atoms in a gas phase can be accumulated in the sample cell when its surfaces are covered by a superfluid helium film. The operation of the cryogenic dissociator which we will describe in the following is based on this effect.
The second dissociation method is based on the natural radioactivity of T. Collins et. al. Collins et al. (1992) first observed that large atomic concentrations accumulate in solid film containing T2 at low temperatures. The dissociation of molecules is done by the -decay electrons generated inside the films. Even a few concentration of T2 or HT in the sample is sufficient to create samples with large concentrations of H and D atoms in various matrices Sheludiakov (2017).
A detailed study of tritiated solid hydrogen films below 1 K was performed at the University of Turku with direct atom diagnostics using magnetic resonance methods: ESR and ENDOR Sheludiakov (2017); Sheludiakov et al. (2017). T2 films of 35 and 250 nm thickness were deposited onto a surface of quartz microbalance (QM) which provided accurate control of the film thickness during deposition and measurements. The QM gold electrode also served as a mirror of the Fabry-Perot resonator connected to an ESR spectrometer operating at 128 GHz. For a film of 250 nm thick, ESR lines of atomic tritium were detected a few minutes after deposition and grew rapidly, reaching a maximum concentration of cm-3 ( relative to the density of molecules) after three hours. Then, depending on storage temperature, periodic heat spikes were observed in the sample cell followed by a 30-40 % decrease in the density of T and decrease in the thickness of the film. This behavior was explained by an explosive recombination of the atoms in the films after they reached some critical density. The explosions were not observed in a thinner T2 film of 35 nm thickness and were suppressed by condensing a superfluid helium film on top of the 250 nm T2 film. The maximum T densities reached during storage decreased by a factor of 1.8 after heating the 250 nm film from 0.1 to 1 K, and the time between explosions has increased by a factor of . Clearly, thermal explosions can be avoided by improving the heat removal from the T2 film.
@@ III.1 Hydrogen and hfs of deuterium
The cryogenic hydrogen dissociator of H operating below 1 K was first realized in the group of W. Hardy at the University of British Columbia Jochemsen et al. (1982); Statt et al. (1985). The method is based on a cryogenic discharge in helium vapor and is similar to the technique mentioned above and used to create H atoms inside solid molecular H2 matrices. The molecules in solid are split by impacts of the electrons of plasma discharge created during short pulses. Part of atoms evaporate into the bulk of the dissociator chamber and are then pushed by magnetic field gradients out of the dissociator. This technique also allowed to create sources of low field seeking H atoms for loading magnetic traps. The transport of atoms to the trap could be done via a short tube, filling the trap with hot gas of atoms with subsequent thermalization and isolation from the walls of the trapHess et al. (1987); Doyle (1991); van Roijen et al. (1988) or following several stages of thermalization and cooling them to mK before entering the trap. The latter technique was described in a previous publication of the Turku group Semakin et al. (2025), where the operation of the cryogenic dissociator for H is described in detail. Atomic fluxes close to s-1 of lfs H are obtained by this technique and were successfully used to load a large magnetic trapSemakin et al. (2025).
@@
It is useful to consider what happens with the helium film and its vapor in the cryogenic dissociator during and after the RF discharge pulse. In a typical operation with H described in a previous work of the Turku group Semakin et al. (2025), the discharge is driven by RF pulses of ms length followed by ms OFF time. The RF power is adjusted to maximum, which the dilution refrigerator can tolerate but is lower than the threshold for full evaporation of the helium film inside the dissociator chamber. The power mW released in the dissociator resonator during the 1 ms RF pulse is sufficient to evaporate moles of 4He which has J/mole latent heat of evaporation at 0.6 K Donnelly . This is nearly half of the total amount of helium that covers the walls of the dissociator. During the operation of the dissociator its average temperature increases from 0.6 to K, which leads to an increase in the 4He vapor pressure in the chamber from cm-3 to cm-3 van Dijk et al. (1960); Kaufman et al. (1993). We used here the average temperatures during pulsed operation measured by a thermometer at the dissociator body. Obviously, the temperature during the pulse is somewhat higher, and therefore the estimate above represents a lower bound for the density of helium vapor during the pulse.
@@
We can estimate the upper limit if we assume that all evaporated during the pulse amount of liquid is instantly converted into vapor. This gives a vapor density of cm-3, quite close to the lower bound estimate. However, this vapor during and after the pulse flushes out of the dissociator and is re-condensed in the colder regions of the transfer line. The vapor is dense enough to entrain D(T) atoms that were created during the discharge pulse, and it pushes them towards the colder regions of the transfer line. This effect, similar to the operation of old diffusion pump is useful for increasing efficiency of the atomic source. It was first observed in the first experiments on stabilization of H Silvera and Walraven (1980a) and received an acronym HEVAC (Helium Vapor Compressor). The superfluid flow along the walls of the transfer line returns helium into the dissociator chamber. Such circulation of helium exists in any system where there is a gradient of temperature and often creates problems because of the associated heat load to the lower temperature parts of the system.
@@
A cryogenic dissociator of this type was also used in experiments with atomic hfs of D, for the hyperfine resonance study of an unpolarized gas in zero field Reynolds et al. (1991); Hayden and Hardy (1995) and with two-dimensional D in a strong magnetic field Mosk et al. (2001). The fluxes were a factor of 5-10 smaller than for H. Attempts to load lfs to a magnetic trap at MIT were unsuccessful Steinberger (2004). The reason may be in the transfer line between the dissociator and the trapping cell, which was very short. The trap was filled with very warm gas, which recombined at subsequent cooling via thermalization with the walls. Clearly, some improvements are needed for the D source for loading magnetic traps with lfs, which we consider in the next section.
@@ III.2 Deuterium and tritium lfs
Our proposal for the cryogenic source of lfs D and T is based on the dissociation technique below 1 K described above. However, further cooling of atoms and their transport to the magnetic trap should not rely on thermalization of atoms by collisions with the superfluid helium covered wall of the transfer line. As mentioned above, the latter may work somehow with hfs of D, may not work well with lfs of D, and highly likely will not work with T. However, the presence of helium cannot be avoided in the dissociator chamber. Helium vapor is necessary for running the discharge in the dissociator. All hydrogen isotopes have a saturated vapor pressure too low to maintain discharge below 1 K. Therefore, we consider the same construction for the D(T) dissociator: copper chamber with a pre-condensed solid layer of D2(T2) on its inner walls and a saturated helium film on top of that. A helical resonator with a resonance frequency in the 100-500 MHz range is located inside; see Fig.1.
@@
The dissociator chamber is located at the edge of a superconductive magnet with a maximum field of 4-5 T. The magnetic field gradient separates the atoms with opposite electron spins, pushing lfs out into a transfer line. We suggest building the transfer line with two sections and thermalization points in between, each at slightly lower temperature (see Fig. 1). A system of sextupole or octupole race-track coils with 6 or 8 linear conductors around each of the transfer line sections creates a radial magnetic field gradient. The fields of 3 T are technically available with modern superconducting technologies. This creates a radial potential barrier of K. For spatially separating the lfs atoms from the wall, their thermal energies need to be about 5-10 times lower than the height of the potential barrier, i.e., the atoms leaving the dissociator should be cooled to 0.2-0.4 K before they experience a collision with the wall. The temperature of the dissociator should be lowered as close as possible to this limit, and the geometry of the transfer line just after it should be optimized for a minimum number of wall collisions.
@@
Lowering the dissociator temperature leads to a very steep decrease in saturated vapor of helium in the considered temperature range. In Fig. 2, we present the dependence of the saturated vapor density of 4He and 3He as a function of temperature van Dijk et al. (1960); Kaufman et al. (1993); Huang and Chen (2006). Since we aim at thermalizing the T gas with the helium vapor, we should have a mean-free path for the collisions of T with He comparable to the diameter of the transfer line, which we assume to be cm. With the T-4He collisional cross-section taken from Table 2, we find that the temperatures should be mK using pure 4He. We can use 3He-4He mixture films to increase the vapor density at a given temperature. Pure 3He cannot be used because the film will not be superfluid. Mixtures up to 5 of 3He were previously used by the Turku cryogenic dissociator of hfs of H Safonov et al. (2001). Taking the saturated vapor density for this mixture, we evaluate the minimum dissociator temperature of mK to have the T gas well thermalized in the 3He vapor. These temperature estimates match the above-mentioned estimate for reliable magnetic isolation from walls with the 3 T magnets. In the following, two modes of operation for the dissociator are considered: with pure 4He (T=0.4 K) and with 5 3He-4He mixture (T=0.2 K) films.
@@
A small temperature gradient in the transfer line will lead to condensation of helium vapor in the colder regions of the transfer line. The vapor pressure gradient will move the T-He mixture towards the colder region until the vapor density vanishes and the D(T) gas gets decoupled from the walls. For the 5 3He-4He mixture film, this will happen somewhere in the middle of the second section and the gas will end up cooled to 170 mK or slightly below. With pure 4He this will occur already in the beginning of the first section and the gas will pass further without cooling.
@@
The KATRIN++ and Project8 projects plan to use large magnetic traps for storage of T atoms for neutrino mass measurements. In principle, the temperature range 0.2-0.4 K achieved with the T source described above is already sufficiently low for magnetic trapping. A superconductive trap with 3.1 T (2 K) magnetic barrier has been built in the High Energy Accelerator Research Organization (KEK) Laboratory in Japan Huffman et al. (2014) and used for trapping ultra-cold neutrons O’Shaughnessy et al. (2009). However, for the large trap volumes considered (up to 10 m3), smaller fields and lower temperatures are preferable. Also, as we discussed in Section IID, the minimal atom loss due to dipolar relaxations is reached at around 0.1 K. In order to increase the efficiency of the buffer gas cooling even under conditions when the mean-free path of T-He collisions is larger than the transfer line diameter, we suggest slowing down the flow of the T beam along the line. This can be done by placing extra pinch coils after each section, which will create potential barriers for the atoms. These coils will create an impedance for gas flow along the transfer line and increase the number of collisions with the low-density buffer gas. Another option to cool the D(T) gas flowing in the transfer line without buffer gas was recently proposed by the Project 8 collaboration Esfahani et al. (2025). They suggest utilizing evaporative cooling, allowing for the high-temperature tail of the energy distribution of the flowing gas to overcome the radial magnetic barrier and stick to the wall.
@@
For the experiments aiming at precision spectroscopy and the observation of the gravitational quantum states with D or T, we suggest feeding the atomic beam out of the cryostat with the dilution refrigerator to room temperature instruments, so that the experiments can be basically done using the same techniques as is done with the 6 K nozzle Matveev et al. (2013); Killian et al. (2024). Since the superfluid helium film is supposed to line the inner walls of the transfer line, extending it to room temperature requires suppression of the film and preferably also He vapor flow together with the atomic beam. This can be done by installing a superfluid film cutter based on the evaporation and condensation of the film in a specially designed geometry Kaufman et al. (1993); Ishikawa et al. (2010) or on a suppression of superfluid flow on the Cs coated surface Anthony-Petersen et al. (2024).
@@
For experiments with the slow beams, we may employ the pulsed mode of operation of the dissociator. The average flux of atoms/s considered above consists of pulses 1 ms long followed by a 20 ms delay. The number of atoms in each pulse is and close to what was reported by Helffrich Helffrich et al. (1987) for a similar type dissociator for H. The beams of D(T) at 0.2-0.4 K can be slowed at room temperature using Zeeman decelerators. This technique is based on applying decelerating magnetic field pulses synchronized with the motion of the bunches of atoms in the beam. For H, a device consisting of 12 decelerating stages decreased atom velocities from 520 to 100 m/s Hogan et al. (2008) with an average deceleration of 35 m/s per stage. Two to three such stages will be enough to completely stop the 0.2-0.4 K cold D (T) atoms exiting from the source proposed in this work.

### v2

@@ II.3 Scattering of H isotopes on each other and on He
Elastic collisions that do not change the spin state of colliding atoms lead to the recovery of thermal equilibrium and are essential for evaporative cooling. Inelastic rates typically lead to a change of the hyperfine state which may lead to the loss of the atoms from magnetic trap as discussed in the next section. Scattering of atoms with helium vapor is important for the transport of hydrogen isotopes into the trapping cell and defines the efficiency of buffer gas cooling, which is one of the key effects in this proposal.
@@
In the zero temperature limit, the elastic collision cross-section is related to the s-wave scattering length as , where is the two-body correlator defined by quantum statistics and identical properties of colliding particles. 1, 2, 0 for two distinguishable particles, for two indistinguishable bosons, and for two identical fermions accordingly. The value and sign of the scattering length depend on the interaction potential and the electron spin state of the atoms of the pair. The singlet potential supports a large number of bound molecular states and is not relevant for spin-polarized or spinless atoms considered here. The repulsive triplet potential has a weak well because of the Van-der-Waals attraction and may support weakly bound states, or dimers. The shape of the potential is nearly the same for all atomic pairs considered here. The reduced mass of the atomic pair turns out to be an important parameter in the scattering properties. The dimer bound state occurs for the heaviest pair considered here 4He-4He pair. The pair of T atoms is very close to the binding resonance [44]. At the scattering resonance, the scattering length diverges and changes sign, going from to . The scattering length and the collisional cross-section then depend on how far the reduced mass is from the resonance value of [44]. Scattering lengths and cross-sections for elastic collisions of various pairs of H and He isotopes in the zero temperature limit were calculated by several authors [45, 37, 46, 47]. We present a summary of the data in Table 2. Of particular significance for this work is a very large elastic collision cross-section for the heaviest pairs considered, and especially for the T-T and T-He collisions [44, 47].
| D-D [46, 45] | T-T [44] | H-T [46] | T-3He [47] | T-4He [47] | 3He-3He [46, 45] | 3He-4He [46, 45] | 4He-4He [46, 45] | |
| , Å | -4 | -42 | -0.85 | -8.8 | -22 | -7 | -17 | -110 |
| Å2 | 2 | 450 | 0.09 | 9.7 | 61 | 6.4 | 36 | 3200 |
Processes that occur in inelastic collisions are the spin-exchange and dipolar relaxation. The first class concerns the collisions of atoms in different or mixed hyperfine states (c for H/T and and for D, see Fig. 1). This process is very important during loading and at the initial stage of magnetic trapping. The exchange and dipolar relaxation rates for numerous channels for atomic hydrogen were calculated in Refs. [48, 44] and experimentally confirmed in the work of the MIT [49, 50] and Amsterdam groups [51].
@@
The situation is very different for fermionic D. Due to the Pauli principle, identical D atoms avoid approaching each other, and s-wave scattering is forbidden. For D, theory predicts two orders of magnitude faster spin-exchange rates, and the dipolar relaxation decreasing linearly with temperature as cm3/s K [52]. Remarkably, the doubly polarized D gas is becoming more stable at lower temperatures, and its decay time at 1 mK is two orders of magnitude larger than that for H, while the thermalization rate remains fairly large [52]. For tritium, relaxation rate constant data were recently calculated [44]. Spin-exchange and dipolar relaxation are much larger than for H, and the rates have a complex dependence on temperature and magnetic field. Under optimal conditions of magnetic fields below 0.01-0.02 T and a temperature of 100 mK the dipolar relaxation rate is 5 times higher than for H. Unfortunately, this implies that the fluxes of T in loading the trap should be increased by the same factor.
II.4 Magnetic trapping of H, D, and T
Since surface recombination limits the cooling of H states below 100 mK, a wall-free confinement method was proposed by Pritchard [53] and later extended to hydrogen by Hess [54]. It is based on trapping low field seekers () in a potential well created by a magnetic field. Cooling of the trapped gas is performed by evaporation of the atoms, which have enough energy to overcome the trapping barrier and carry away the energy. The remaining gas is thermalized in elastic two-body collisions. Successive trapping and cooling experiments were performed at MIT [49] and the University of Amsterdam [51] reaching BEC in 1998 [4]. Recently, a large magnetic trap for H was built and tested in the University of Turku [55]. The typical range of H densities confined in magnetic traps is between and cm-3 and minimum temperatures are several tens of K.
The trappable low field seeking states c (, in zero field basis) and d (, ) are higher in energy than two other hyperfine states a (, ) and b (, ) (see Fig.1). Relaxation to the lower energy states is threshold-less and does not vanish even at zero temperature. Spin exchange during two-body collisions is the fastest loss channel. It occurs in collisions where the ”mixed” state c takes part leading to a rapid depletion of the c state in the trap. The remaining ”pure” and doubly polarized state d is much more stable since spin-exchange does not work in this case. However, the dipolar interaction during collision of two d atoms may lead to spin flip and relaxation to untrapped states. This channel of dipolar relaxation is the main loss mechanism that limits the lifetime and maximum densities of the trapped gas. This is a second order, in density, process with the characteristic decay time inversely proportional to the gas density. For H, the two-body dipolar relaxation rate constant is cm3/s [48], which implies a lifetime of s at the density cm-3.
Stability of the trapped T gas is basically determined by relaxation processes similar to those in H. Both atoms are bosons and have a similar hyperfine structure. However, the three fold larger mass has a strong effect. As recently calculated [44], the low-energy cross sections for T-T collisions are substantially larger than those for H-H. This concerns both elastic and inelastic collisions. The reason is that a scattering resonance appears in collisions of atoms with a hypothetical mass slightly greater than 3 [44]. Two T atoms are very close to the formation of the bound dimer state. The s-wave scattering length is very large and negative Å. This increases the elastic collision rate and enhances the efficiency of evaporative cooling. However, inelastic dipole relaxation is increased as well. For the relaxation d-d of two T atoms in a magnetic field of 1 mT and temperature of 1 mK, typical for H trapping, the increase reaches a factor of compared to that for H-H collisions. The dipolar rate for T-T collisions has a much stronger temperature dependence, and the difference with H decreases to a factor of for T=100 mK. This leads to an increase in the loss rate and a decrease in the life-time of the trapped T gas by the same factors at equal density. To our knowledge, magnetic trapping of T has not been attempted to date, and the theoretical predictions above are still waiting for experimental confirmation.
As mentioned in the previous section, trapping of atomic D would allow substantially larger densities and very high stability of the trapped gas [52]. Compared with the potential of all three isotopes for precision spectroscopy in the magnetic traps, deuterium is obviously the best candidate. However, the only attempt to trap the D gas performed by the MIT group using the same dissociation and trapping techniques as for H [31] was unsuccessful. The walls of their trapping cell were covered with superfluid helium film. Although some flux of D entering the trap was detected, no signal was seen after the dissociator was switched off. The large adsorption energy on the helium film and fast recombination rate on the surface of the trapping cell were proposed as a possible explanation for their failure to load the trap.
@@ II.5 Stabilization of H, D, and T in solid molecular matrices
Several methods are known to produce the matrices with embedded atoms. We briefly describe two methods relevant to this work: dissociation with cryogenic RF discharge and electrons resulting from the -decay of T2. The first technique is based on an in-situ dissociation of the matrix molecules using pulsed RF discharge in the helium vapor above solid films of hydrogen below 1 K introduced by Ahokas et. al. [56]. The molecules are split by the impact of the electrons of the discharge having energies of the order of 100 eV. It turned out that a fraction of the atoms is evaporated during the RF pulse, and the H and D atoms in a gas phase can be accumulated in the sample cell when its surfaces are covered by a superfluid helium film. The operation of the cryogenic dissociator which we will describe in the following is based on this effect.
The second dissociation method is based on the natural radioactivity of T, which is mixed into solid films [57, 58]. The dissociation of molecules is caused by the -decay electrons generated inside the films. Even a few concentration of T2 or HT in the sample is sufficient to produce samples with large ( cm-3) concentrations of H and D atoms in various matrices [59].
A detailed study of tritiated solid hydrogen films below 1 K was performed at the University of Turku with direct atom diagnostics using magnetic resonance methods: ESR and ENDOR [59, 60]. T2 films of 35 and 250 nm thickness were deposited onto the surface of a quartz microbalance (QM) which provided accurate control of the film thickness during deposition and measurements. The QM gold electrode also served as a mirror of the Fabry-Perot resonator connected to an ESR spectrometer operating at 128 GHz. For a film of 250 nm thick, ESR lines of T were detected a few minutes after deposition and grew rapidly, reaching a maximum concentration of cm-3 ( relative to the density of molecules) after three hours. Then, depending on the storage temperature, periodic heat spikes were observed in the sample cell followed by a 30-40 % decrease in the density of T and decrease in the thickness of the film. This behavior was explained by an explosive recombination of the atoms in the films after they reached some critical density. The explosions were not observed in a thinner T2 film of 35 nm thickness and were suppressed by condensing a superfluid helium film on top of the 250 nm T2 film. The maximum T densities reached during storage decreased by a factor of 1.8 after heating the 250 nm film from 0.1 to 1 K, and the time between explosions has increased by a factor of . Clearly, thermal explosions can be avoided by improving the heat removal from the T2 film or if the atoms can be removed from the solid matrix (e.g. by running the RF discharge) before a large concentration is accumulated.
@@ III.1 Hydrogen and hfs of deuterium
The cryogenic hydrogen dissociator of H operating below 1 K was first realized in the group of W. Hardy at the University of British Columbia [61, 62]. The method is based on a cryogenic discharge in a helium vapor and is similar to the technique mentioned above and used to produce H atoms inside solid molecular H2 matrices. The molecules in solid are split by impacts of the electrons of plasma discharge produced during short pulses. Part of atoms evaporate into the bulk of the dissociator chamber and are then pushed by magnetic field gradients out of the dissociator. This technique also allowed to design sources of low field seeking H atoms for loading magnetic traps. The transport of atoms to the trap could be done via a short tube, filling the trap with hot gas of atoms with subsequent thermalization and isolation from the walls of the trap [49, 50, 51] or following several stages of thermalization and cooling them to mK before entering the trap. The latter technique was described in a previous publication of the Turku group [25], where the operation of the cryogenic dissociator for H is described in detail. Atomic fluxes close to s-1 of lfs H are obtained by this technique and were successfully used to load a large magnetic trap [25].
@@
It is useful to consider what happens with the helium film and its vapor in the cryogenic dissociator during and after the RF discharge pulse. In a typical operation with H [25], the discharge is driven by RF pulses of ms length followed by ms OFF time. The RF power is adjusted to the maximum value, which the dilution refrigerator can tolerate but is lower than the threshold for full evaporation of the helium film inside the dissociator chamber. The energy released in the dissociator resonator during the 1 ms RF pulse is sufficient to evaporate moles of 4He which has J/mole latent heat of evaporation at 0.6 K [63]. This is nearly half of the total amount of helium that covers the walls of the dissociator. During the dissociator operation, its average temperature increases from 0.6 to K, leading to an increase in the 4He vapor density in the chamber from cm-3 to cm-3 [64, 65]. We used here the average temperatures during pulsed operation measured by a thermometer at the dissociator body. Obviously, the temperature during the pulse is somewhat higher, and therefore the estimate above represents a lower bound for the density of helium vapor during the pulse.
@@
We can estimate the upper limit if we assume that the amount of liquid evaporated during the pulse is instantly converted into vapor. This gives a vapor density of cm-3, quite close to the lower bound estimate. However, this vapor during and after the pulse flushes out of the dissociator and is re-condensed in the colder regions of the transfer line. The vapor is dense enough to entrain D(T) atoms that were produced during the discharge pulse, and it pushes them towards the colder regions of the transfer line. This effect, similar to the operation of the old diffusion pump, is useful to increase the efficiency of the atomic source. It was first observed in the first experiments on the stabilization of H [2] and received the acronym HEVAC (Helium Vapor Compressor). The superfluid flow along the walls of the transfer line returns helium into the dissociator chamber. Such circulation of helium exists in any system where there is a gradient of temperature and often creates problems because of the associated heat load to the lower temperature parts of the system.
@@
A cryogenic dissociator of this type was also used in experiments with atomic hfs of D, for the hyperfine resonance study of an unpolarized gas in a zero field [41, 27] and with two-dimensional D in a strong magnetic field [36]. The fluxes were a factor of 5-10 smaller than for H. Attempts to load the lfs to a magnetic trap at MIT were unsuccessful [31]. The reason may be in the transfer line between the dissociator and the trapping cell, which was very short. The trap was filled with very warm gas, which recombined at subsequent cooling via thermalization with the walls. Clearly, some improvements are needed for the D source for loading magnetic traps with lfs, which we consider in the next section.
@@ III.2 Deuterium and tritium lfs
Our proposal for the cryogenic source of lfs D and T is based on the dissociation technique below 1 K described above. However, further cooling of atoms and their transport to the magnetic trap should not rely on thermalization of atoms by collisions with the superfluid helium covered wall of the transfer line. As mentioned above, the latter may work somehow with hfs of D, may not work well with lfs of D, and highly likely will not work with T. However, the presence of helium cannot be avoided in the dissociator chamber. Helium vapor is necessary for running the discharge in the dissociator. All hydrogen isotopes have too low saturated vapor pressures to maintain discharge below 1 K. Therefore, we consider the same construction for the D(T) dissociator: copper chamber with a pre-condensed solid layer of D2(T2) on its inner walls and a saturated helium film on top of that. A helical resonator with a resonance frequency in the 100-500 MHz range is located inside; see Fig. 2.
@@
The dissociator chamber is located at the edge of a superconductive magnet with a maximum field of 4-5 T. The magnetic field gradient separates the atoms with opposite electron spins, pushing lfs out into a transfer line. We propose building the transfer line with two sections and thermalization points between, each at slightly lower temperature (see Fig. 2). A system of sextupole or octupole race-track coils with 6 or 8 linear conductors around each of the transfer line sections provides a radial magnetic field gradient. The fields of 3 T are technically available with modern superconducting technologies for such race-track type of coils. This provides a radial potential barrier of K. For effective isolation of the lfs atoms from the wall, their thermal energies must be 5-10 times lower than the height of the potential barrier, i.e., the atoms leaving the dissociator should be cooled to 0.2-0.4 K. The temperature of the dissociator should be lowered as close as possible to this limit, and the geometry of the transfer line just after it should be optimized for a minimum number of wall collisions.
@@
Lowering the dissociator temperature leads to a very steep decrease in the saturated vapor density of helium in the considered temperature range. In Fig. 3, we present the dependence of the saturated vapor density of 4He and 3He as a function of temperature [66, 67]. Since we aim to thermalize the T gas with the helium vapor, we should have a mean-free path for the collisions of T with He comparable to the diameter of the transfer line, which we assume to be cm. Using the temperature-dependent cross sections for the collisions of T with 4He from ref. [47] we find that the mean-free path for T-4He collisions becomes equal to 1 cm at 440 mK.
We can use 3He-4He mixture films to increase the vapor density at a given temperature. Pure 3He cannot be used because the film will not be superfluid. Mixtures up to of 3He were previously used by the Turku cryogenic dissociator of hfs of H [33]. It is known that the vapor pressure above a regular solution of 3He in 4He is higher than , the product of saturated vapor pressure at the same temperature and the concentration of 3He in the liquid [68]. The concentration of 3He in the film inside a tube with a temperature gradient should increase towards the warmer end due to an osmotic effect [68]. Considering that both the above-mentioned effects lead to an enhancement of the 3He density in the dissociator, we take as a lower limit estimate for the vapor density above the mixture. Then, we evaluate the minimum dissociator temperature mK for which the T mean free-path in the 3He vapor becomes equal to 1 cm. These temperature estimates match the above-mentioned value for reliable magnetic isolation from walls using 3 T magnets. In the following, two modes of operation for the dissociator are considered: with pure 4He (T=0.44 K) and with 5 3He-4He mixture (T=0.22 K) films.
@@
A small temperature gradient in the transfer line will lead to condensation of helium vapor in the colder regions of the transfer line. The vapor pressure gradient will move the T-He mixture towards the colder region until the vapor density vanishes and the D(T) gas gets decoupled from the walls. For the 5 3He-4He mixture film, this will happen somewhere in the middle of the second section and the gas will end up cooled to 220 mK or slightly below. With pure 4He this will occur already in the beginning of the first section and the gas will pass further without cooling.
Efficiency of the buffer gas cooling can be increased by increasing the length of the atom guide after the dissociator. In the estimates above we set a constraint on the mean-free path of 1 cm. However, the length of the section following the dissociator can be as large as 1 m. Exchange of energy in the elastic collisions of T with helium is very efficient due to the closeness of masses. Each collision reduces the energy difference by the factor of two, and 2-3 collisions are sufficient for good thermalization [26]. In this case, reaching the low temperature end of the atom guiding line, T gas will undergo several collisions at temperatures below 0.2(0.4) K.
@@
The KATRIN++ and Project8 projects plan to use large magnetic traps for storage of T atoms for neutrino mass measurements. In principle, the temperature range of 0.22-0.44 K achieved with the T source described above is already sufficiently low for magnetic trapping. A superconductive trap with 3.1 T (2 K) magnetic barrier has been built in the High Energy Accelerator Research Organization (KEK) Laboratory in Japan [69] and used to trap ultra-cold neutrons [70]. However, for the large trap volumes considered (up to 10 m3), smaller fields and lower temperatures are preferable. Also, as we discussed in Section II D, the minimal atom loss due to dipolar relaxation is reached at around 0.1 K. In order to increase the efficiency of the buffer gas cooling even under the condition that the mean-free path of T-He collisions is larger than the transfer line diameter, we propose slowing down the flow of the T beam along the line. This can be done by placing extra pinch coils after each section, which will provide potential barriers for the atoms. These coils will provide an impedance for gas flow along the transfer line and increase the number of collisions with the low-density buffer gas. Another option to cool the D(T) gas flowing in the transfer line without buffer gas was recently proposed by the Project 8 collaboration [71, 72]. They propose using evaporative cooling that allows the high-temperature tail of the energy distribution of the flowing gas to overcome the radial magnetic barrier and be removed.
@@
For experiments aiming at precision spectroscopy and the observation of the gravitational quantum states of D or T, we propose feeding the atomic beam out of the cryostat with the dilution refrigerator to room temperature instruments, so that the experiments can be done using the same techniques as is done with a 6 K nozzle [73, 9]. Since the superfluid helium film is supposed to line the inner walls of the transfer line, extending it to room temperature requires suppression of the film and preferably also He vapor flow together with the atomic beam. This can be done by installing a superfluid film cutter based on the evaporation and condensation of the film in a specially designed geometry [65, 74] or by suppressing superfluid flow on the surface coated with Cs [75].
@@
For experiments with slow beams, we may employ the pulsed mode of operation of the dissociator. The average flux of atoms/s considered above consists of pulses 1 ms long followed by a delay of 20 ms. The number of atoms in each pulse is and close to that reported by Helffrich [76] for a similar type dissociator for H. The D(T) beams at 0.2-0.4 K can be slowed at room temperature using Zeeman decelerators. This technique is based on applying decelerating magnetic field pulses synchronized with the motion of the bunches of atoms in the beam. For H, a device consisting of 12 deceleration stages decreased atom velocities from 520 to 100 m/s [77] with an average deceleration of 35 m/s per stage. Two to three such stages will be enough to completely stop the 0.2-0.4 K cold D (T) atoms exiting from the source proposed in this work.